OpenStudy (anonymous):

which ordered pair is the solution to this equation? -2x+9y=-26 a.(-4,-4) b.(4,4) c.(-4,-5) d.(-5,-4)

OpenStudy (anonymous):

plug in x and y of each and see if you get -26

OpenStudy (anonymous):

wait can you explain it more

myininaya (myininaya):

for example choice a is the ordered pair (x=-4,y=-4) -2(-4)+9(-4)=8-36=-28 which is not -26 so choice a is not an ordered pair on the line -2x+9y=-26

myininaya (myininaya):

try d (x=-5,y=-4) -2(-5)+9(-4)=10-36=-26 which equals -26 so (-5,-4) is on the line -2x+9y=-26 d is the correct solution